The velocity of particle moving in straight line is decreasing at rate of 3m/s per metre of displacement at an instant when the velocity is 10m/s. Determine the acceleration of the particle at this instant.

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear Chilukuri

given dv/ds=3

or v=3s

 or a =dv/dt =3 ds/dt

                  =3v  

         a= 3*10 =30 m/sec2


Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the

answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin